Solve the system
2x-4y-2z=12
x-y+z=7
-5x+5y-3z=-33

Answer by kekegeter (97)


2x-4y-2z=12
2x=4y+2z+12
x=2y+2+6
___________
-5x+5y-3z=-33
-5x=-5y+3z-33
x=- -5y+3z-33/5
_______________
x-y+z=7
x=y-2+7